# Beyond the Neural Fog: Interpretable Learning for AC Optimal Power Flow

### v1

@@ I Introduction
The traditional approach to solving the OPF problem involves a DC approximation, which is formulated as a linear programming problem with a global optimal solution [2]. This method has been widely adopted and underpins many market-clearing algorithms. However, its limitations in capturing the complexities of power systems are becoming increasingly apparent. While the DC approximation offers computational efficiency, it often leads to significant economic losses and suboptimal solutions. More importanly, the solution provided by the DC approximation is always AC-infeasible [3]. The growing prevalence of reactive power and voltage constraints in contemporary networks further exacerbates the accuracy concerns associated with this approach [4]. The technical literature also proposes other linearized versions of the OPF problem beyond the traditional DC-OPF. For instance, [5] introduces a linearization of OPF based on loss factors. Similarly, [6] presents a bilevel problem to find the best parameters to approximate OPF with linear equations.
@@
More recently, machine learning techniques have been introduced to solve the OPF problem [8]. Deep neural networks have been employed to learn the active set of constraints in AC-OPF [9], while sensitivity-informed neural networks have been proposed to address the AC-OPF problem [10]. Furthermore, graph neural networks have emerged as a potential solution for tackling the AC-OPF challenge [11]. These methods offer rapid computation, facilitating real-time implementation. However, they typically require extensive datasets and are challenging to train. Moreover, the black-box nature of neural networks raises concerns about transparency and interpretability, making system operators hesitant to deploy them in real networks.
In this paper, we propose a learning-based methodology to reformulate the OPF problem into a linear optimization problem that closely approximates the AC-OPF solution. Our approach combines model- and learning-based methodologies by utilizing a first-order Taylor expansion of the original AC power flow equations around an estimated operating point derived from historical data. Similar to other model-based approximations, a global optimal solution to the linear programming model we propose can be achieved using standard optimization algorithms. However, unlike the well-known DC-OPF approximation, our approach adapts to each operating point, resulting in solutions that are closer to those of the original AC-OPF problem. Similar to other learning-based approaches, ours leverages historical data to improve accuracy. However, unlike approaches based on neural networks, our method is easy to train, transparent, and interpretable, making it more suitable for adoption in real-world power systems.
Figure 1 visually compares the modus operandi of neural networks and our approach. Neural networks (NN) seek to learn a complex nonlinear function that fits all observed data points, represented as black dots. In contrast, our method (KNN+Taylor) focuses on a localized linear approximation. By employing a nearest neighbor clustering algorithm, we identify the region of interest centered around the anticipated optimal power flow solution. Within this region, we apply a Taylor expansion to the original power flow equations to construct a linear model.
@@ II Formulations
Consider a power network composed of nodes, branches connecting the nodes, and generators. The nodes, indexed by $n\in\mathcal{N}$, are characterized by a shunt conductance ($G^{sh}_{n}$), a shunt susceptance ($B^{sh}_{n}$), an active power demand ($p^{d}_{n}$), and a reactive power demand ($q^{d}_{n}$). The voltage magnitude and angle at each node are denoted by $v_{n}$ and $\theta_{n}$, respectively, and the fomer is bounded by technical limits $\underline{v}_{n}$, $\overline{v}_{n}$. The generators, indexed by $g\in\mathcal{G}$, are characterized by a convex cost function $c_{g}()$ and minimum/maximum active and reactive power outputs denoted by $\underline{p}_{g}/\overline{p}_{g}$ and $\underline{q}_{g}/\overline{q}_{g}$, respectively. The branches, indexed by $l\in\mathcal{L}$, are characterized by a series resistance ($r_{l}$), a series reactance ($x_{l}$), a total charging susceptance ($b_{l}$), a tap ratio magnitude ($\tau_{l}$), and a phase shift angle ($\gamma_{l}$). We define two sparse connection matrices $F$ and $T$ of size ||×$|\mathcal{L}|\times|\mathcal{N}|$ such that $F_{ln}$ and $T_{lm}$ are equal to 1 if node $n/m$ is the start/end node of branch $l$, with all other elements being zero. The active power flow from/to each branch $l$ are denoted by $p^{f}_{l}$ and $p^{t}_{l}$, respectively, while $q^{f}_{l}$ and $q^{t}_{l}$ similarly represent the reactive power flows. For each branch the apparent power flow is limited by $\overline{s}_{l}$ and we also define:
@@
Model (3) is convex and therefore can be solved to global optimality with off-the-shelf optimization software. As discussed in [13], the DC formulation of the OPF offers several advantages. Firstly, its solutions are non-iterative, reliable, and unique, ensuring consistency and dependability in results. Secondly, the methods and software used in this formulation are relatively simple and user-friendly. Thirdly, the DC-OPF can be solved and optimized efficiently, which is particularly beneficial in the demanding area of contingency analysis. Additionally, it requires minimal network data that is relatively easy to obtain. The linear nature of the DC-OPF aligns well with the economic theories on which the majority of current transmission-constrained electricity markets are based. Lastly, the approximated MW flows provided by the DC-OPF are reasonably accurate, especially for the heavily loaded branches that might constrain system operation.
@@
where $\tilde{p}^{f}_{l},\tilde{q}^{f}_{l},\tilde{p}^{t}_{l},\tilde{q}^{t}_{l}$ are the power flows of the given operating point obtained through equations (2d)–(2g). Despite involving a greater number of variables and constraints, optimization model (4) retains the same structural advantages as model (3). Its convex objective function and constraints ensure that it can be solved to global optimality in a finite number of steps using efficient algorithms. However, unlike the DC-OPF formulation, model (4) incorporates both reactive power flows and voltage magnitudes, which allows reducing the degree of infeasibility and/or the cost of repair of the so-obtained approximate solutions. The primary challenge with this method lies in selecting an appropriate operating point for the Taylor expansion. To address this, we introduce in Section III a simple, fast, and interpretable learning-based algorithm to infer this operating point, specifically utilizing a $K$-nearest neighbor regression technique [19].
@@ III Learning-based methods
In this section, we describe the various learning-based approaches compared in this paper to infer the solution of the AC-OPF problem using nodal active and reactive demands as features. Let us denote $\mathbf{d}=[(p^{d}_{n},q^{d}_{n})\forall n]$ as the vector containing the active and reactive demands at all nodes. We characterize the solution of the optimal power flow by the active power output of the generators $\mathbf{p}=[(p_{g})\forall g]$ and the voltage magnitudes at the nodes with generating units $\mathbf{v}=[(v_{n})\forall n\in\mathcal{N}_{G}]$, where $\mathcal{N}_{G}$ is the subset of nodes with generating units. Using $\mathbf{d},\mathbf{p},\mathbf{v}$, we can solve a conventional AC power flow to obtain the remaining variables, including reactive generation levels, voltage magnitudes at load buses, voltage angles in the network, and active and reactive power flows through all branches. The learning methods analyzed in this paper assume the availability of a dataset containing historical solutions of the AC-OPF model (2) denoted by ($\mathbf{p}^{*},\mathbf{v}^{*}$) for different demand profiles ($\mathbf{d}$). Specifically, we have access to $\{(\mathbf{d}_{i},\mathbf{p}^{*}_{i},\mathbf{v}^{*}_{i}),\forall i\in\mathcal{I}\}$. We use this training data to infer as accurately as possible the solution of the AC-OPF for a new instance $i^{\prime}$ with a demand profile $\mathbf{d}_{i^{\prime}}$.
@@
As discussed in Section I, finding the global optimal solution for AC-OPF instances is computationally demanding due to the problem’s non-convex nature and the absence of convergence guarantees. However, we assume that the historical solutions used in our analysis have been computed offline by solving model (2) to global optimality, without constraints on computational resources. This assumption aligns with the standard practice in the literature for learning-based methods addressing the AC-OPF problem.
@@
First, we consider a $K-$nearest neighbor (KNN) regression that infers the solution to the AC-OPF problem as the average of the solutions of the nearest neighbors according to a given distance [19]. In particular, for a given value of $K$, we determine the subset of instances $\mathcal{I}_{K}$, with $|\mathcal{I}_{K}|=K$, that have the lowest Euclidean distance with respect to $\mathbf{d}_{i^{\prime}}$. Then, the solution is inferred as the average of the closest neighbors, that is, $\hat{\mathbf{p}}_{i^{\prime}}=(\sum_{i\in\mathcal{I}_{K}}\mathbf{p}^{*}_{i})/K$ and $\hat{\mathbf{v}}_{i^{\prime}}=(\sum_{i\in\mathcal{I}_{K}}\mathbf{v}^{*}_{i})/K$. This is a simple and interpretable learning method that can achieve good accuracy, especially with large training datasets. However, it is purely data-driven and lacks inherent knowledge of the underlying physics governing the problem. This can lead to performance degradation when the input data deviates significantly from the training set, as the closest neighbors may not accurately represent the new instance $i^{\prime}$.
@@
At the other end of the spectrum lies neural networks (NN), which boast high versatility and have demonstrated their ability to capture complex, non-linear relationships between inputs and outputs in optimization problems due to their vast number of parameters [20]. However, this very complexity comes at a cost. Neural networks are often criticized for their lack of interpretability, making it difficult to understand how they arrive at their solutions. Additionally, they require substantial training data and can be challenging to train effectively. In this paper, we employ two distinct neural networks to predict the active power output of generators ($\mathbf{p}$) and the voltage magnitudes at the buses with generation ($\mathbf{v}$). Detailed information about the neural network architecture used for learning these variables of the AC-OPF solution is provided in Section V.
@@
The learning approach proposed in this paper begins with a KNN regression to estimate an operating point close to the AC-OPF solution for a given demand profile, $\mathbf{d}_{i^{\prime}}$. This operating point is then used to solve an AC power flow problem, where the active and reactive demand values ($\mathbf{d}_{i^{\prime}}$) are fixed. The active power generation for all units, except the slack bus, is set to the values estimated by the KNN ($\hat{\mathbf{p}}_{i^{\prime}}$), and the voltage magnitudes at the generation nodes are also fixed to the KNN estimates ($\hat{\mathbf{v}}_{i^{\prime}}$). Although the operating point derived from solving the power flow problem might violate certain inequality constraints (2h)-(2l), it inherently satisfies the power balance and power flow equality constraints (2b)-(2g) of the AC-OPF problem. This property makes it a suitable reference for calculating a first-order Taylor expansion of the nonlinear power flow equations (2d)–(2g) around this point. The final step involves solving the linearized models (4) or (5) to determine the solution of the linearized optimal power flow problem. This hybrid method combines the strengths of model-based and learning-based approaches, offering a balance between interpretability, computational efficiency, and the ability to capture the physical relationships inherent in the power flow problem. Two variants of the proposed approach are presented: KT-AC and KT-DC. The KT-AC variant solves the linear problem (4), which accounts for both active and reactive power flows through the branches. In contrast, the KT-DC variant solves the linear problem (5), focusing solely on the relationship between active power flow and voltage angles, while neglecting reactive power and voltage magnitudes.
@@
Occasionally, due to the Taylor expansion, optimization problems (4) or (5) may become infeasible. When this occurs, the solution reverts to the one provided by the KNN approach. This should not be confused with the general fact that the solutions derived from all model- and learning-based approximation methods (DC-OPF, KNN, NN, KT-AC, KT-DC) are generally not AC-feasible, as they often fail to satisfy some constraints of the original AC-OPF model (2).

### v2

@@ I Introduction
The traditional approach to solving the OPF problem involves a DC approximation, which is formulated as a linear programming problem with a global optimal solution [2]. This method has been widely adopted and underpins many market-clearing algorithms. However, its limitations in capturing the complexities of power systems are becoming increasingly apparent. While the DC approximation offers computational efficiency, it often leads to significant economic losses and suboptimal solutions. More importantly, the solution provided by the DC approximation is always AC-infeasible [3]. The growing prevalence of reactive power and voltage constraints in contemporary networks further exacerbates the accuracy concerns associated with this approach [4]. The technical literature also proposes other linearized versions of the OPF problem beyond the traditional DC-OPF. For instance, [5] introduces a linearization of OPF based on loss factors. Similarly, [6] presents a bilevel problem to find the best parameters to approximate OPF with linear equations.
@@
More recently, machine learning techniques have been introduced to solve the OPF problem [8]. Deep neural networks have been employed to directly learn the solution of the DC-OPF [9] or to learn the active set of constraints in AC-OPF [10], while sensitivity-informed, physics-informed and graph neural networks have been proposed to address the AC-OPF problem [11, 12, 13, 14]. These methods offer rapid computation, facilitating real-time implementation. However, they typically require extensive datasets and are challenging to train. Moreover, the black-box nature of neural networks raises concerns about transparency and interpretability, making system operators hesitant to deploy them in real networks.
In this paper, we propose a learning-based methodology to reformulate the OPF problem into a convex optimization problem that closely approximates the AC-OPF solution. Our approach combines model- and learning-based methodologies by utilizing a first-order Taylor expansion of the original AC power flow equations around an estimated operating point derived from historical data. Similar to other model-based approximations, a global optimal solution to the convex programming model we propose can be achieved using standard optimization algorithms. However, unlike the well-known DC-OPF approximation, our approach adapts to each operating point, resulting in solutions that are closer to those of the original AC-OPF problem. Similar to other learning-based approaches, ours leverages historical data to improve accuracy. However, unlike approaches based on neural networks, our method is easy to train, transparent, and interpretable, making it more suitable for adoption in real-world power systems.
Figure 1 visually compares the modus operandi of neural networks and our approach. Neural networks (NN) seek to learn a complex nonlinear function that fits all observed data points, represented as black dots. In contrast, our method (KNN+Taylor) focuses on a localized linear approximation: By employing a $K$-nearest neighbor (KNN) clustering algorithm, we identify the region of interest centered around the anticipated optimal power flow solution and, within this region, we apply a Taylor expansion to linearly approximate the power flow equations. Since the neural network aims to approximate the function across its entire domain, it requires a large amount of training data and is challenging to train. In contrast, the proposed approach requires significantly less data by approximating the function locally and dynamically adapting it to each specific instance.
@@ II Formulations
Consider a power network composed of nodes, branches connecting the nodes, and generators. The nodes, indexed by $n\in\mathcal{N}$, are characterized by a shunt conductance ($G^{sh}_{n}$), a shunt susceptance ($B^{sh}_{n}$), an active power demand ($p^{d}_{n}$), and a reactive power demand ($q^{d}_{n}$). Given the inherent complexity of the optimal power flow in its deterministic formulation, we assume throughout this work that electricity demand is known with certainty. Addressing the stochastic version of this problem is considered beyond the scope of this study. The voltage magnitude and angle at each node are denoted by $v_{n}$ and $\theta_{n}$, respectively, and the fomer is bounded by technical limits $\underline{v}_{n}$, $\overline{v}_{n}$. The generators, indexed by $g\in\mathcal{G}$, are characterized by a convex cost function $c_{g}()$ and minimum/maximum active and reactive power outputs denoted by $\underline{p}_{g}/\overline{p}_{g}$ and $\underline{q}_{g}/\overline{q}_{g}$, respectively. The branches, indexed by $l\in\mathcal{L}$, are characterized by a series resistance ($r_{l}$), a series reactance ($x_{l}$), a total charging susceptance ($b_{l}$), a tap ratio magnitude ($\tau_{l}$), and a phase shift angle ($\gamma_{l}$). We assume all branch parameters are known. We define two sparse connection matrices $F$ and $T$ of size ||×$|\mathcal{L}|\times|\mathcal{N}|$ such that $F_{ln}$ and $T_{lm}$ are equal to 1 if node $n/m$ is the start/end node of branch $l$, with all other elements being zero. The active power flow from/to each branch $l$ are denoted by $p^{f}_{l}$ and $p^{t}_{l}$, respectively, while $q^{f}_{l}$ and $q^{t}_{l}$ similarly represent the reactive power flows. For each branch the apparent power flow is limited by $\overline{s}_{l}$ and we also define:
@@
Model (3) is convex and therefore can be solved to global optimality with off-the-shelf optimization software. As discussed in [16], the DC formulation of the OPF offers several advantages. Firstly, its solutions are non-iterative, reliable, and unique, ensuring consistency and dependability in results. Secondly, the methods and software used in this formulation are relatively simple and user-friendly. Thirdly, the DC-OPF can be solved and optimized efficiently, which is particularly beneficial in the demanding area of contingency analysis. Additionally, it requires minimal network data that is relatively easy to obtain. The convex nature of the DC-OPF aligns well with the economic theories on which the majority of current transmission-constrained electricity markets are based. Lastly, the approximated MW flows provided by the DC-OPF are reasonably accurate, especially for the heavily loaded branches that might constrain system operation.
@@
where $\tilde{p}^{f}_{l},\tilde{q}^{f}_{l},\tilde{p}^{t}_{l},\tilde{q}^{t}_{l}$ are the power flows of the given operating point obtained through equations (2d)–(2g). Despite involving a greater number of variables and constraints, optimization model (4) retains the same structural advantages as model (3). Its convex objective function and constraints ensure that it can be solved to global optimality in a finite number of steps using efficient algorithms. However, unlike the DC-OPF formulation, model (4) incorporates both reactive power flows and voltage magnitudes, which allows reducing the degree of infeasibility and/or the cost of repair of the so-obtained approximate solutions. Although model (4) is a convex approximation of the true model (2), it retains a similar level of interpretability, as it provides insights into the binding physical constraints and explains why further reductions in dispatch costs are not possible without violating the technical limits of the generating units or network branches. This represents a significant advantage over black-box methods based on neural networks, which rely on complex, non-transparent mappings learned from data, making it challenging to understand how the results are derived or to trace the reasoning behind the solutions.
Although Taylor expansion has previously been used to iteratively solve the AC-OPF problem [22, 23, 24], this is, to the best of our knowledge, the first time that learning techniques have been employed to compute, based on historical data, the Taylor expansion around an operating point expected to be close to the optimal solution of the AC-OPF. Furthermore, unlike other procedures that use Taylor approximations, our approach is non-iterative, which significantly reduces the overall computational burden. Since the first-order Taylor expansion is more accurate near the operating point where it is applied, the solution of model (4) approximates that of model (2) as long as its optimal solution is close enough to the operating point used for the Taylor expansion. Therefore, the primary challenge with this method lies in selecting an appropriate operating point for the Taylor expansion. To address this, we introduce in Section III a simple, fast, and intuitive learning-based algorithm to infer this operating point, specifically utilizing a $K$-nearest neighbor regression technique [25].
@@ III Learning-based methods
In this section, we describe the various learning-based approaches compared in this paper to infer the solution of the AC-OPF problem using nodal active and reactive demands as features. Let us denote $\mathbf{d}=[(p^{d}_{n},q^{d}_{n})\forall n]$ as the vector containing the active and reactive demands at all nodes. We characterize the solution of the optimal power flow by the active power output of the generators $\mathbf{p}=[(p_{g})\forall g]$ and the voltage magnitudes at the nodes with generating units $\mathbf{v}=[(v_{n})\forall n\in\mathcal{N}_{G}]$, where $\mathcal{N}_{G}$ is the subset of nodes with generating units. Using $\mathbf{d},\mathbf{p},\mathbf{v}$, we can solve a conventional AC power flow to obtain the remaining variables, including reactive generation levels, voltage magnitudes at load buses, voltage angles in the network, and active and reactive power flows through all branches.
@@
As discussed in Section I, finding the global optimal solution for AC-OPF instances is computationally demanding due to the problem’s non-convex nature and the absence of convergence guarantees. However, we assume that the historical solutions used in our analysis have been computed offline by solving model (2) to global optimality, without constraints on computational resources. In this context, and aligned with common assumptions in the supervised learning literature, this work presumes the availability of a dataset containing historical globally optimal solutions to the AC-OPF model (2) denoted by ($\mathbf{p}^{*},\mathbf{v}^{*}$) for different demand profiles ($\mathbf{d}$). Specifically, we have access to $\{(\mathbf{d}_{i},\mathbf{p}^{*}_{i},\mathbf{v}^{*}_{i}),\forall i\in\mathcal{I}\}$, where $\mathcal{I}$ is the set of training instances. In practice, when optimal solutions for large networks are unavailable, state estimation outputs can serve as proxies, as these operating points statistically satisfy physical constraints and have near-minimal costs.
@@
We use this training data to infer as accurately as possible the solution of the AC-OPF for a new instance $i^{\prime}$ with a demand profile $\mathbf{d}_{i^{\prime}}$. First, we consider a $K-$nearest neighbor (KNN) regression that infers the solution to the AC-OPF problem as the average of the solutions of the nearest neighbors according to a given distance [25]. In particular, for a given value of $K$, we determine the subset of instances $\mathcal{I}_{K}$, with $|\mathcal{I}_{K}|=K$, that have the lowest Euclidean distance with respect to $\mathbf{d}_{i^{\prime}}$. Then, the solution is inferred as the average of the closest neighbors, that is, $\hat{\mathbf{p}}_{i^{\prime}}=(\sum_{i\in\mathcal{I}_{K}}\mathbf{p}^{*}_{i})/K$ and $\hat{\mathbf{v}}_{i^{\prime}}=(\sum_{i\in\mathcal{I}_{K}}\mathbf{v}^{*}_{i})/K$. This is a simple and intuitive learning method that can achieve good accuracy, especially with large training datasets. However, it is purely data-driven and lacks inherent knowledge of the underlying physics governing the problem. This can lead to performance degradation when the input data deviates significantly from the training set, as the closest neighbors may not accurately represent the new instance $i^{\prime}$.
@@
At the other end of the spectrum of learning-based approaches are neural networks (NN), which boast high versatility and have demonstrated their ability to capture complex, non-linear relationships between inputs and outputs in optimization problems due to their vast number of parameters [26]. However, this very complexity comes at a cost. Neural networks are often criticized for their lack of interpretability, making it difficult to understand how they arrive at their solutions. Additionally, they require substantial training data and can be challenging to train effectively. In this paper, we employ two distinct neural networks to predict the active power output of generators ($\mathbf{p}$) and the voltage magnitudes at the buses with generation ($\mathbf{v}$). Detailed information about the neural network architecture used for learning these variables of the AC-OPF solution is provided in Section V.
@@
The learning approach proposed in this paper begins with a KNN regression to estimate an operating point close to the AC-OPF solution for a given demand profile, $\mathbf{d}_{i^{\prime}}$. This operating point is then used to solve an AC power flow problem, where the active and reactive demand values ($\mathbf{d}_{i^{\prime}}$) are fixed. The active power generation for all units, except the slack bus, is set to the values estimated by the KNN ($\hat{\mathbf{p}}_{i^{\prime}}$), and the voltage magnitudes at the generation nodes are also fixed to the KNN estimates ($\hat{\mathbf{v}}_{i^{\prime}}$). Although the operating point derived from solving the power flow problem might violate certain inequality constraints (2h)-(2l), it inherently satisfies the power balance and power flow equality constraints (2b)-(2g) of the AC-OPF problem. This property makes it a suitable reference for calculating a first-order Taylor expansion of the nonlinear power flow equations (2d)–(2g) around this point. The final step involves solving the convex models (4) or (5) to determine the solution of the optimal power flow problem. This hybrid method combines the strengths of model-based and learning-based approaches, offering a balance between interpretability, computational efficiency, and the ability to capture the physical relationships inherent in the power flow problem. Two variants of the proposed approach are presented: KT-AC and KT-DC. The KT-AC variant solves the convex problem (4), which accounts for both active and reactive power flows through the branches. In contrast, the KT-DC variant solves the convex problem (5), focusing solely on the relationship between active power flow and voltage angles, while neglecting reactive power and voltage magnitudes. Algorithm 1 includes a pseudo-code of the proposed approach.
Algorithm 1 Pseudo-code of proposed approach
Input: Historical data $\{(\mathbf{d}_{i},\mathbf{p}^{*}_{i},\mathbf{v}^{*}_{i}),\forall i\in\mathcal{I}\}$, number of neighbors $K$, and test demand profile $\mathbf{d}_{i^{\prime}}$.
Output: Active power output of generators $\mathbf{p}_{i^{\prime}}$ and voltage magnitudes at generating buses $\mathbf{v}_{i^{\prime}}$.
1:  Determine $K$ closest neighbors $\mathcal{I}_{K}$.
2:  Compute $\hat{\mathbf{p}}_{i^{\prime}}=\frac{\sum_{i\in\mathcal{I}_{K}}\mathbf{p}^{*}_{i}}{K}$ and $\hat{\mathbf{v}}_{i^{\prime}}=\frac{\sum_{i\in\mathcal{I}_{K}}\mathbf{v}^{*}_{i}}{K}$.
3:  Solve power flow with $\mathbf{d}_{i^{\prime}}$,$\hat{\mathbf{p}}_{i^{\prime}}$,$\hat{\mathbf{v}}_{i^{\prime}}$ to obtain $\{\tilde{v}_{n},\tilde{\theta}_{n},\forall n\}$.
4:  Compute derivatives (7) and solve (4) or (5).
5:  If feasible, return $\mathbf{p}_{i^{\prime}},\mathbf{v}_{i^{\prime}}$. Otherwise, return $\hat{\mathbf{p}}_{i^{\prime}},\hat{\mathbf{v}}_{i^{\prime}}$.
@@
Although the methods mentioned above do not constitute an exhaustive list of all approaches in the literature aimed at solving the AC-OPF problem, we believe they represent a broad spectrum of methodologies, including classical and widely used model-based approximations (e.g., DC-OPF), simple and interpretable learning-based methods (e.g., KNN), and more complex black-box learning methods (e.g., neural networks).
At this point, we must make an important clarification. For any given demand $\mathbf{d}_{i^{\prime}}$, the DC-OPF, KNN, and NN approaches always provide a corresponding operating point. However, optimization problems (4) or (5) may occasionally become infeasible for certain demands $\mathbf{d}_{i^{\prime}}$ and the operating point $\{\tilde{v}_{n},\tilde{\theta}_{n},\forall n\}$ used for the Taylor expansion. In such rare cases, one can address infeasibilities by solving a modified version of problems (4) and (5), incorporating slack variables to increase the maximum power flow capacities of network branches and adjust the maximum/minimum voltage levels at buses. These slack variables are penalized in the objective function, ensuring that they take non-zero values only when the original models are infeasible. However, given the extremely low number of infeasible instances and in the interest of simplicity, we have chosen in this paper to directly utilize the solutions provided by the KNN approach in those few cases, as they offer similar levels of accuracy to those achieved by the approach with slack variables.
This should not be confused with the general observation that solutions derived from all model- and learning-based approximation methods (DC-OPF, KNN, NN, KT-AC, KT-DC) are typically not AC-feasible, as they often fail to satisfy certain constraints of the original AC-OPF model (2).